OpenStudy (anonymous):

Bohrium-267 has a half-life of 15 seconds. Find the amount that remains from a 16 mg sample of this substance after 2 minutes. f(x)=p*a^x

OpenStudy (anonymous):

Easy when dealing with multiples of half-lift times. 120 s / 15 s = 8 fraction remaining is (1/2)^8 = 1/256 amount is this fraction times starting amount.

OpenStudy (anonymous):

huh . lol im confussed

OpenStudy (tkhunny):

15 seconds 1/2 30 seconds 1/4 45 seconds 1/8 60 seconds 1/16 75 seconds 1/32 90 seconds 1/64 ... Getting the idea?

OpenStudy (anonymous):

Lol not really .

OpenStudy (anonymous):

Good explanation.

OpenStudy (anonymous):

 i have to fill that in .